Given the following information provide an analysis that answers the questions at the end of the assignment.

Purchase Price: $4,100,000

Rents:

Tenant

Unit 1

Unit 2

Unit 3

Unit 4

Unit 5

Unit 6

Unit 7

Unit 8

Monthly Rent

$5,600

$5,600

$5,800

$3,200

$3,200

$3,200

$3,200

$3,200

Year 1 NOI: $259,578

NOI Growth Rate: 3%

Loan:

               LTV: 75%

               Amortization Period: 27 years, Term: 10 years

               Interest Rate: 4.35%

               Lender Points: 2%

Selling Costs: 5%

Going Out Cap Rate: 6.5%

Discount Rate: 11%

Questions:

  1. What is the going in cap rate?
  2. What is year 1 PGI?
  3. What is the DSCR?
  4. What is your Initial Investment?
  5. What is the sale price after the five-year hold?
  6. What is the loan balance at sale?
  7. What is the NPV for this project?
  8. What id the IRR for this project?
  9. What is the DCF for this project?
  10. What is the reversion amount?

Excel formulas please
Answer #1

1. Going in cap rate = NOI/Purchase price = 259,578/41,00,000 = 6.33%

2. Year 1 PGI = 33,000*12= 3,96,000

3 DSCR

Net Operating Income (NOI) = Gross Operating Income - Vacancy Loss - Operating Expenses

NOI = $396,000 - $0 - $136,422

NOI = $259,578

Debt Service = Loan Amount * Interest Rate / 100 / [1 - (1 + Interest Rate / 100 / 12) (-12 * Loan Term) )]

Debt Service = $3,075,000 * 4.35 / 100 / [1 - (1 + 4.35 / 100 / 12) (-12 * 10) )]

Debt Service = $379,763.24

Debt Service Coverage Ratio (DSCR) = NOI / Debt Service

DSCR = $259,578 / $379,763.24

DSCR = 0.68 x

4. Initial investment = Purchase price * Loan % = 41,00,000*75% = 30,75,000

5. Sale price after 5 year hold :

Amortisation amount for 5 years =41,00,000/27*5 = 759,259

Amortised value = 41,00,000 - 759,259 = 3340740

Value after selling cost of 5% = 3340740 - (41,00,000*5%) = 3135740

6. IRR = Initial investment = 30,75,000

Cash flows = 396000

IRR = -33.42%